### v3

# GenCeption: Evaluate Multimodal LLMs with Unlabeled Unimodal Data

Abstract
Multimodal Large Language Models (MLLMs) are typically assessed using expensive annotated multimodal benchmarks, which often lag behind the rapidly evolving demands of MLLM evaluation. This paper outlines and validates GenCeption, a novel, annotation-free evaluation method that requires only unimodal data to measure inter-modality semantic coherence and inversely assesses MLLMs’ tendency to hallucinate. This approach eliminates the need for costly data annotation, minimizes the risk of training data contamination, results in slower benchmark saturation, and avoids the illusion of emerging abilities. Inspired by the DrawCeption game, GenCeption begins with a non-textual sample and proceeds through iterative description and generation steps. The semantic drift across iterations is quantified using the GC@$T$ metric. Based on the GenCeption method, we establish the MMECeption benchmark for evaluating Vision LLMs (VLLMs), and compare performance of several popular VLLMs and human annotators. Our empirical results validate GenCeption’s effectiveness, demonstrating strong correlations with established VLLM benchmarks. VLLMs still significantly lack behind human performance and struggle especially with text-intensive tasks.
GenCeption: Evaluate Multimodal LLMs with Unlabeled Unimodal Data
[…]
1 Introduction
Large Language Models (LLMs) demonstrate exceptional abilities in natural language understanding, reasoning, and problem-solving. Multimodal LLMs (MLLMs) enhance these capabilities by incorporating multiple modalities, with the visual modality being predominant Achiam et al. (2023); Liu et al. (2023b); Jiang et al. (2023); Ye et al. (2023). Building on LLMs, MLLMs integrate non-textual modalities, enabling richer interactions and broader applications in real-world scenarios. However, there is a lack of comprehensive evaluation methods to compare different MLLM architectures and training approaches Fu et al. (2023).
In response, the community has developed several MLLM benchmarks, as detailed by Xu et al. (2022); Dai et al. (2023); Wang et al. (2023); Ye et al. (2023); Li et al. (2023b); Zhao et al. (2023). Yet, these benchmarks face common challenges: (1) They predominantly rely on multimodal datasets that demand high-quality annotations, which is costly and restrictive in capturing the evolving capabilities of MLLMs Fu et al. (2023). This has been shown to result in increasing speed in benchmark saturation while contemporary models still struggle on trivial real-world tasks Kiela et al. (2021). (2) MLLM evaluation benchmarks rely on discrete metrics like accuracy that may falsely suggest emergent abilities and do not allow predictable projections of performance improvements from model scaling Schaeffer et al. (2023). (3) The evaluation scores may not reflect true performance on real-world tasks due to potential contamination of MLLM training data by benchmark datasets, as reported for LLM pretraining corpora Dodge et al. (2021); Yang et al. (2023). (4) The content of one modality is often not needed to answer benchmark questions, as the answer can often be inferred from the question or the MLLM’s pretraining knowledge. As a consequence of both (3) and (4), some MLLMs can excel on vision QA benchmarks without even being provided the image that is associated with the question.
Figure 1: An illustration of the $t$-th iteration in the GenCeption evaluation procedure for MLLMs. Using the image modality as an example, the process begins with an existing image $\mathbf{X}^{(0)}$ sourced from a unimodal image dataset for the first iteration ($t$=1). The MLLM provides a detailed description of the image, which is then used by an image generator to produce $\mathbf{X}^{(t)}$.
To address these highlighted challenges, we propose GenCeption, a novel and simple approach for evaluating MLLMs. GenCeption addresses challenge (1) by relying on easily accessible unimodal datasets, which allows for cost-effective and scalable benchmark creation. Relying on unimodal datasets additionally addresses challenge (3) and (4), as it allows to easily use previously unseen datasets for MLLM evaluation, and enforces the relevance of the provided modality for excelling at this task. To tackle challenge (2), GenCeption uses the continuous GC@$T$ metric, providing a more nuanced evaluation compared to discrete metrics, allowing for better projections of performance improvements and avoiding false suggestions of emergent abilities.
GenCeption assesses MLLMs’ ability to consistently maintain semantic coherence across modalities by iteratively generating and describing non-textual samples and measures the continuous GC@$T$ metric. This approach simultaneously evaluates the MLLM’s tendency to hallucinate, as this inversely correlates with semantic coherence. Further, an MLLM’s ability to provide complete yet concise descriptions of non-textual samples measures a diverse range of specialised abilities. For instance, to perform well at describing an image using a limited number of tokens, it is advantageous to be able to reason over people’s emotions and intentions behind their actions, infer the current and preceding weather, count objects, and recognize artistic styles. This list can be extended to various abilities depending on the non-textual modality and the content of the samples used during the GenCeption process.
Our contributions are the following:
  • Proposing GenCeption, an evaluation method that allows for using unlabelled unimodal datasets for MLLM evaluation.
  • Introducing MMECeption, a Vision LLM (VLLM) evaluation benchmark utilizing the GenCeption method. MMECeption uses the images from the MME benchmark Fu et al. (2023), but without their annotated question and answer pairs.
  • Evaluating seven leading VLLMs on the MMECeption benchmark and comparing results with other popular VLLM benchmarks and human performance.
We will elaborate on the proposed implementation of the GenCeption method, detail our experimental setup, and discuss our findings.
2 GenCeption
Our approach, GenCeption, is inspired by a multi-player game DrawCeption [^1] https://wikipedia.org/wiki/drawception (a.k.a., Scrawl or Whispernary). In this game, the first player is given an image and describes it verbally to the next player. This player then attempts to recreate the image based on the description. The cycle repeats, often resulting in amusing deviations from the original image. The challenge and objective are to maintain the initial information through iterative transitions between verbal descriptions and drawings. Similarly, a proficient Multimodal Language Model (MLLM), which models multiple modalities such as text and images, should excel at this task. Recognizing that MLLMs can encompass modalities beyond just visual cues, such as audio and graphs, we name our approach GenCeption, covering a broader scope than the visually-centric DrawCeption.
While it may not be possible to preserve the initial information perfectly due to varying levels of richness, accuracy, and ambiguity in different modalities, a more capable MLLM will minimize the semantic drift from the original input. This contrasts with common benchmarks that aim for complete saturation, highlighting a key advantage of the GenCeption framework: the creation of benchmarks that are more challenging to saturate. With complex initial samples, such as images of real-world scenes or graphs with numerous nodes and edges, this may even result in impossible-to-saturate benchmarks. Aiming for minimum rather than no semantic drift, this would allow to rank MLLMs relative to each other while continuously leaving space for more performant models.
2.1 Procedure
Unlike existing MLLM benchmarks that rely on multimodal samples, GenCeption is designed to operate on unimodal datasets, significantly streamlining dataset acquisition efforts. For illustrative purposes, we employ the image modality as a representative non-textual modality throughout this exposition. Let’s consider an image dataset $\bm{\mathcal{D}}$ comprising images $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{N}$, similar to well-established datasets like ImageNet Deng et al. (2009), CIFAR Krizhevsky et al. (2009), and STL Coates et al. (2011). Without loss of generality, any image from $\bm{\mathcal{D}}$ is denoted as $\mathbf{X}$.
Please write a clear, precise, detailed, and concise description of all elements in the image. Focus on accurately depicting various aspects, including but not limited to the colors, shapes, positions, styles, texts and the relationships between different objects and subjects in the image. Your description should be thorough enough to guide a professional in recreating this image solely based on your textual representation. Remember, only include descriptive texts that directly pertain to the contents of the image. You must complete the description using less than 500 words.
Table 1: The fixed textual prompt $\mathbf{P}_{\text{Desc}}$ instructs the MLLM to produce a description of the input $\mathbf{X}^{(t-1)}$.
GenCeption operates iteratively, from $t=1$ to a pre-defined maximum iteration $t=T$. Each iteration, as depicted in Figure 1, begins with an image $\mathbf{X}^{(t-1)}$ and yields a new image $\mathbf{X}^{(t)}$. The first iteration ($t=1$) starts with the original image $\mathbf{X}^{(0)}$ from $\bm{\mathcal{D}}$. During any given iteration $t$, the MLLM receives a textual prompt $\mathbf{P}_{\text{Desc}}$ (Table 1), instructing the MLLM to generate a comprehensive description $\mathbf{Q}_{t}$ for the input image $\mathbf{X}^{(t-1)}$:
$\mathbf{Q}_{t}:=\text{MLLM}(\mathbf{P}_{\text{Desc}},\mathbf{X}^{(t-1)}).$ (1)
Following this, an image generation prompt $\mathbf{P}_{\text{Gen}}^{(t)}$ is constructed as “Generate an image that fully and precisely reflects this description: <$\mathbf{Q}_{t}$>”. This prompt guides a pretrained image generation model, such as DALL·E Ramesh et al. (2021) or Imagen DeepMind (2023), to create a new image, $\mathbf{X}^{(t)}$:
$\mathbf{X}^{(t)}:=\text{Gen}(\mathbf{P}_{\text{Gen}}^{(t)}),$ (2)
where Gen(·) signifies the chosen image generator. Each subsequent iteration $t+1$ starts with the image $\mathbf{X}^{(t)}$ generated in the previous iteration. Upon completing all iterations, we obtain a series of $T+1$ images: $\mathbf{X}^{(0)},\mathbf{X}^{(1)},\ldots,\mathbf{X}^{(T)}$, with the initial image being the original and the rest sequentially produced across the iterations.
The textual prompt $\mathbf{P}_{\text{Desc}}$ is intentionally kept short and concise to minimize potential variations in model behaviours due to susceptibility to prompt composition Loya et al. (2023).
2.2 Metric: GC@$T$
Our primary objective is to measure the semantic divergence of each generated image $\mathbf{X}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$) from the original image $\mathbf{X}^{(0)}$. We use a pretrained image encoder, such as ViT Dosovitskiy et al. (2021), to transform all images, resulting in $T+1$ image embeddings denoted as $\mathbf{z}^{(0)},\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(T)}$, where $\mathbf{z}^{(t)}:=\text{Enc}(\mathbf{X}^{(t)})$. We then compute the cosine similarity between $\mathbf{z}^{(0)}$ and each $\mathbf{z}^{(t)}$ (for $t\in\{1,2,\ldots,T\}$), yielding $T$ similarity scores: $s^{(1)},s^{(2)},\ldots,s^{(T)}$. Here, $s^{(t)}\in[-1.0,1.0]$ approximates the level of semantic drift in the $t$-th iteration of the GenCeption procedure. To quantify the overall speed and magnitude of semantic drift, we calculate the GenCeption score over $T$ iterations, denoted as GC@$T\in[-1.0,1.0]$, as follows:
[…]
This is a normalized and continuous metric that weights later iterations more heavily for two reasons: (1) similar to the DrawCeption game, the last image’s deviation from the initial image is most telling; (2) we aim to capture performance and dynamics across the entire iterative sequence. A high GC@$T$ value signifies an exceptional ability to maintain inter-modal (text-image) semantic congruence, effectively curbing the propensity for rapid or extensive deviation from the semantics encapsulated in the original image. Notably, GC@$1$ is equivalent to $s^{(1)}$. For the pseudo code detailing the GenCeption procedure and the calculation of the average GC@$T$ metric over the entire dataset $\bm{\mathcal{D}}$, see Algorithm 1.
[…]
$\text{GC}_{\text{FID}}\text{@}T:=\frac{\sum_{t=1}^{T}(t\cdot\text{fid}^{(t)})}{\sum_{t=1}^{T}t}.$ (4)
[…]
Input: MLLM to be evaluated, a unimodal dataset $\bm{\mathcal{D}}$:
$\smash{\mathbf{X}_{1}^{(0)},\ldots,\mathbf{X}_{n}^{(0)},\ldots,\mathbf{X}_{N}^{(0)}}$, fixed textual prompt $\mathbf{P}_{\text{Desc}}$,
a sample generator Gen(·), and a sample encoder Enc(·)
Output: Average GC@$T$ metric over $\bm{\mathcal{D}}$
1
Parameter: The number of iterations $T$
2
3GC@$T$ = 0
4for $(n=1;n\leq N;n++)$ do
5
6      $\mathbf{z}^{(0)}$ := Enc($\mathbf{X}_{n}^{(0)}$);
7      for $(t=1;t\leq T;t++)$ do
8
9            Generate description $\mathbf{Q}_{t}$ for $\mathbf{X}_{n}^{(t-1)}$ using (1);
10            Create sample generation prompt $\smash{\mathbf{P}_{\text{Gen}}^{(t)}}$;
11            Generate a new sample $\mathbf{X}_{n}^{(t)}$ according to (2);
12            $s^{(t)}$ := CosineSimilarity($\smash{\mathbf{z}^{(0)}}$, Enc($\smash{\mathbf{X}_{n}^{(t)}}$));
13       end for
14
15      Calculate GC@$T$ += $\smash{\sum_{t=1}^{T}(t\cdot s^{(t)})/\sum_{t=1}^{T}t}$; (3)
16 end for
17
18return GC@$T$ / N;
Algorithm 1 Calculate GC@$T$ via GenCeption
Sample group & category Gemini1.5-Pro Claude3-Opus GPT-4o GPT-4V mPLUG-Owl2 LLaVA-13B LLaVA-7B
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
[…]
GC@3$\uparrow$
Cosine
GC@3$\downarrow$
FID
visual-intensive samples Existence 190.0 0.437 269.8 183.3 0.382 273.7 195.0 0.400 266.5 175.0 0.422 265.1 185.0 0.323 296.9 195.0 0.305 322.0 195.0 0.308 318.1
Count 148.3 0.389 272.6 116.7 0.348 285.3 190.0 0.388 277.5 153.3 0.404 277.4 160.0 0.299 316.1 165.0 0.294 319.7 148.3 0.253 326.1
Position 105.0 0.357 253.7 76.7 0.357 266.1 145.0 0.398 260.6 85.0 0.408 253.3 75.0 0.306 294.0 135.0 0.255 298.9 123.3 0.285 286.0
Color 175.0 0.474 234.8 118.3 0.385 267.6 180.0 0.421 246.1 141.7 0.403 243.7 138.3 0.290 310.1 165.0 0.300 305.6 170.0 0.284 304.5
Poster 175.2 0.374 206.0 149.7 0.360 206.0 192.2 0.335 203.7 187.8 0.324 209.4 154.8 0.243 209.2 163.6 0.215 240.9 154.1 0.214 244.3
Celebrity 169.4 0.362 191.0 77.6 0.317 192.5 46.8 0.331 193.3 53.5 0.332 189.1 167.9 0.232 211.3 144.4 0.206 223.7 153.2 0.188 233.6
Scene 147.0 0.423 173.7 149.8 0.374 174.7 148.5 0.401 171.5 141.2 0.393 173.7 157.8 0.299 194.4 162.8 0.277 198.0 160.8 0.266 196.3
Landmark 176.8 0.375 182.1 113.0 0.344 188.9 175.5 0.372 182.0 104.0 0.353 182.6 158.8 0.275 206.0 150.8 0.242 224.3 154.8 0.252 214.4
Artwork 152.2 0.412 171.1 136.8 0.385 170.4 144.0 0.415 169.2 115.0 0.421 170.3 136.0 0.252 202.2 98.8 0.212 213.3 110.0 0.210 215.3
Commonsense 150.0 0.464 216.5 115.0 0.432 210.4 174.3 0.448 213.9 155.0 0.471 208.1 127.9 0.353 237.2 115.7 0.334 248.9 117.1 0.294 254.4
Visual Mean 158.9 0.407 217.1 123.7 0.368 223.6 159.1 0.391 218.4 131.2 0.393 217.3 146.2 0.287 247.7 149.7 0.264 259.5 148.7 0.255 259.3
Visual Rank 2 1 1 7 4 4 1 3 3 6 2 2 5 5 5 3 6 7 5 7 6
textual-intensive Code reasoning 117.5 0.213 310.0 70.0 0.245 267.4 182.5 0.255 299.7 147.5 0.193 302.9 65.0 0.176 327.6 55.0 0.144 323.5 50.0 0.107 398.2
Numerical calc. 110.0 0.268 346.5 67.5 0.229 349.3 170.0 0.282 346.4 80.0 0.240 322.5 45.0 0.192 362.0 35.0 0.195 367.4 50.0 0.155 366.0
Text translation 162.5 0.240 334.6 45.0 0.236 362.5 192.5 0.211 326.9 55.0 0.157 368.0 112.5 0.081 365.2 85.0 0.116 352.3 65.0 0.111 424.4
OCR 170.0 0.367 233.2 167.5 0.362 245.5 192.5 0.362 246.2 177.5 0.393 238.0 102.5 0.276 255.4 95.0 0.239 270.6 65.0 0.222 283.7
Textual Mean 140.0 0.272 306.1 87.5 0.268 306.2 184.4 0.278 304.8 115.0 0.246 307.9 81.3 0.181 327.6 67.5 0.174 328.5 57.5 0.149 368.1
Textual Rank 2 2 2 4 3 3 1 1 1 3 4 4 5 5 5 6 6 6 7 7 7
Overall Mean 153.5 0.368 242.5 113.3 0.340 247.2 166.3 0.359 243.1 126.5 0.351 243.2 127.6 0.257 270.5 126.1 0.238 279.2 122.6 0.225 290.1
Overall Rank 2 1 1 7 4 4 1 2 2 4 3 3 3 5 5 5 6 6 6 7 7
HallusionBench* 45.2 (rank=3) 37.8 (rank=4) 51.7 (rank=1) 46.5 (rank=2) 25.7 (rank=6) 24.5 (rank=7) 27.6 (rank=5)
MMStar* 38.6 (rank=5) 45.7 (rank=3) 61.6 (rank=1) 47.7 (rank=2) 34.8 (rank=6) 40.1 (rank=4) 34.6 (rank=7)
SEEDBench (Test)* 70.7 (rank=3) 64.0 (rank=7) 76.4 (rank=1) 71.6 (rank=2) 64.5 (rank=6) 67.9 (rank=4) 66.4 (rank=5)
AI2D* 70.2 (rank=4) 70.6 (rank=3) 82.2 (rank=1) 75.5 (rank=2) 55.7 (rank=7) 61.3 (rank=5) 55.9 (rank=6)
OpenCompass* 62.7 (rank=3) 57.7 (rank=4) 66.3 (rank=1) 63.3 (rank=2) 46.3 (rank=7) 48.8 (rank=5) 46.7 (rank=6)
* Results are sourced from https://huggingface.co/spaces/opencompass/open_vlm_leaderboard as of 2024-04-25 (except GPT-4o) and 2024-05-23 (for GPT-4o).
$\downarrow$ Results are obtained using FID to measure the similarity between images. https://github.com/GaParmar/clean-fid
Table 2: Evaluation results of GC@$3$, MME, HallusionBench and OpenCompass on visual(Vis)-intensive and textual(Text)-intensive images. Best results per metric and category (over different MLLMs) are bolded.
3 Experiments
We run several extensive experiments to validate the GenCeption method by comparing the GC@$T$ scores achieved by several VLLMs to the scores they achieve on carefully crafted established benchmarks and to average human performance. Although GenCeption’s design merely requires unimodal image datasets, we employ the same data as used by a recent and well-validated MLLM benchmark, MME Fu et al. (2023). While we discard the annotated question-answer pairs associated with the images in this benchmark, this provides us with the ability (1) to facilitate direct comparison with metrics that include textual QA (question-answering) annotations, and (2) to enable a detailed assessment of MLLM performance across MME’s 14 meticulously crafted sample categories. Attributing this newly created benchmark to the MME dataset and the GenCeption method, we refer to it as the MMECeption benchmark.
We select seven VLLMs – Gemini1.5-Pro Reid et al. (2024), GPT-4o OpenAI (2024) , GPT-4V Achiam et al. (2023), Claude3-Opus Anthropic (2023), LLaVA-7B/13B Liu et al. (2023b) and mPLUG-Owl2 Ye et al. (2023) – based on their superior performance on the OpenCompass multimodal leaderboard OpenCompass (2023), which incorporates a comprehensive set of benchmarks like MME Fu et al. (2023), HallusionBench Liu et al. (2023a), MMStar Chen et al. (2024), SeedBench Li et al. (2023a), and AI2D Kembhavi et al. (2016). We use DALL·E as the default image generation model. To exclude a bias towards MLLMs developed by OpenAI which may have had access to DALL·E generated images during training, we perform an additional evalutaion of all VLLMs on the GC@$1$ score using Imagen2 as an image generation model. We set the temperature parameter to 0 in both the MLLMs and image generators to minimize the stochasticity in model outputs.
As humans are well versed at integrating vision and language modalities, we aim to quantify average human performance on the MMECeption benchmark. As the GenCeption procedure is a labor-intensive and time-consuming task for humans, we randomly select 5 images from each MME category, and by providing human annotators with the same prompts as defined in Table 1, collect results and calculate the GC@$1$ metric. 5 human annotators (3 master students, 1 lecturer, and 1 artist) were recruited to describe one image of each category such that each image in a category is described by a different person to mitigate personal performance differences. The annotators were given 14 weeks to perform this task and were awarded a generous reimbursement of €40 each to ensure sufficient dedication. All annotators were either native English speakers or fluent at a professional level.
Sample group & category Gemini1.5-Pro Claude3-Opus GPT-4o GPT-4V mPLUG-Owl2 LLaVA-13B LLaVA-7B
Dalle3 Imgn2 Dalle3 Imgn2 Dalle3 Imgn2 Dalle3 Imgn2 Dalle3 Imgn2 Dalle3 Imgn2 Dalle3 Imgn2
visual-intensive samples Existence 0.505 0.529 0.500 0.532 0.536 0.521 0.505 0.530 0.427 0.515 0.416 0.485 0.418 0.506
Count 0.456 0.489 0.466 0.490 0.456 0.494 0.498 0.506 0.378 0.463 0.408 0.466 0.341 0.416
Position 0.511 0.491 0.495 0.480 0.469 0.460 0.501 0.473 0.346 0.452 0.359 0.454 0.350 0.402
Color 0.545 0.525 0.489 0.501 0.480 0.473 0.506 0.490 0.345 0.471 0.420 0.457 0.318 0.436
Poster 0.455 0.388 0.450 0.381 0.445 0.383 0.444 0.365 0.338 0.357 0.303 0.312 0.305 0.266
Celebrity 0.417 0.384 0.424 0.382 0.418 0.373 0.433 0.389 0.319 0.336 0.284 0.317 0.263 0.313
Scene 0.511 0.490 0.504 0.478 0.482 0.474 0.497 0.474 0.385 0.417 0.355 0.404 0.350 0.392
Landmark 0.500 0.485 0.460 0.492 0.494 0.479 0.458 0.480 0.363 0.351 0.376 0.357 0.334 0.333
Artwork 0.494 0.454 0.508 0.461 0.500 0.455 0.504 0.455 0.333 0.385 0.308 0.333 0.294 0.304
Common. 0.545 0.531 0.535 0.507 0.562 0.526 0.563 0.535 0.425 0.493 0.429 0.473 0.417 0.458
Vis Mean 0.494 0.477 0.483 0.470 0.484 0.464 0.491 0.470 0.366 0.424 0.366 0.406 0.339 0.383
Vis Rank 1 1 4 2 3 4 2 2 5 5 5 6 7 7
textual-intensive Code 0.364 0.177 0.304 0.180 0.395 0.179 0.333 0.263 0.281 0.100 0.260 0.168 0.186 0.108
Numerical 0.322 0.417 0.333 0.389 0.366 0.456 0.325 0.383 0.322 0.225 0.336 0.265 0.259 0.222
Text trans. 0.396 0.227 0.356 0.258 0.444 0.277 0.359 0.238 0.173 0.052 0.200 0.118 0.212 0.073
OCR 0.462 0.500 0.486 0.448 0.421 0.441 0.482 0.417 0.358 0.384 0.368 0.385 0.351 0.320
Text Mean 0.386 0.330 0.370 0.319 0.407 0.338 0.375 0.325 0.284 0.190 0.291 0.234 0.252 0.181
Text Rank 2 2 4 4 1 1 3 3 6 6 5 5 7 7
Overall Mean 0.463 0.435 0.451 0.427 0.462 0.428 0.458 0.428 0.343 0.357 0.344 0.357 0.314 0.325
Overall Rank 1 1 4 4 2 2 3 2 6 5 5 5 7 7
Table 3: The impact of different image encoders, DALL·E 3 (Dalle3) vs. Imagen 2 (Imgn2), on GC@$1$ score. Best results per configuration and category (over different MLLMs) are bolded.
3.1 Quantitative results
We partition the 14 MME categories into two groups based on content type: visual-intensive (10 categories: existence, count, position, color, poster, celebrity, scene, landmark, artwork, and commonsense reasoning) and textual-intensive (4 categories: code reasoning, numerical calculation, text translation, and OCR). GC@$3$ scores on the MMECeption benchmark and accuracy on the original MME benchmark are reported per category and as aggregations in Table 2. Additionally, we include the scores and ranks of all evaluated MLLMs on the OpenCompass OpenCompass (2023), MME Fu et al. (2023), HallusionBench Liu et al. (2023a), MMStar Chen et al. (2024), SeedBench Li et al. (2023a), and AI2D Kembhavi et al. (2016) leaderboards. Notably, Gemini1.5-Pro leads our rankings, followed by GPT-4o, GPT-4v, Claude3-Opus, mPLUG-Owl2, and LLaVA-13B/7B. The GenCeption method shows robustness to the similarity metric used, as the overall ranking remains identical when using cosine similarity or FID distance for calculating GC@$T$ scores.
Figure 2: Correlation Matrix of GC@$1$ and GC@$3$ scores on MMECeption, and several other benchmarks.
Sample group
& category
Gemini1.5
-Pro
Claude3
-Opus
GPT-4o GPT-4V
mPLUG
-Owl2
LLaVA
-13B
LLaVA
-7B
Human
$\Delta$between
human & best
visual-intensive samples Existence 0.5841 0.4563 0.5578 0.5434 0.3967 0.3524 0.3782 0.6402 + 9.6045%
Count 0.4140 0.3799 0.2725 0.4882 0.2364 0.3535 0.2038 0.5476 + 12.1671%
Position 0.5546 0.4959 0.4086 0.5639 0.3527 0.4285 0.3899 0.6409 + 13.6549%
Color 0.7081 0.6206 0.6139 0.5516 0.4047 0.4314 0.3506 0.8380 + 18.3449%
Poster 0.5046 0.4362 0.4939 0.4681 0.3998 0.3208 0.2905 0.5456 + 8.1252%
Celebrity 0.4182 0.3988 0.4369 0.4447 0.3714 0.2545 0.2160 0.4671 + 5.0371%
Scene 0.6080 0.5828 0.5229 0.5919 0.4842 0.3906 0.4057 0.6236 + 2.5658%
Landmark 0.4903 0.4932 0.5236 0.5702 0.3613 0.4174 0.3845 0.6045 + 6.0154%
Artwork 0.3725 0.5304 0.5297 0.5252 0.2938 0.2924 0.2336 0.5421 + 2.2059%
Commonsense 0.4338 0.5375 0.5047 0.4012 0.3244 0.4153 0.3532 0.6417 + 19.3860%
Visual Mean 0.5088 0.4932 0.4865 0.5148 0.3625 0.3657 0.3206 0.6091 + 9.7107%
text-intensive Code reasoning 0.3689 0.4085 0.4043 0.3690 0.2923 0.2963 0.1975 0.5376 + 31.6034%
Numerical calc. 0.3652 0.3958 0.3940 0.4241 0.3474 0.4409 0.3423 0.5160 + 17.0333%
Text translation 0.4480 0.3949 0.4333 0.3803 0.0931 0.2372 0.1981 0.6196 + 38.3036%
OCR 0.4382 0.4329 0.3334 0.4455 0.2663 0.3371 0.2912 0.4696 + 5.4097%
Textual Mean 0.4051 0.4080 0.3913 0.4047 0.2498 0.3279 0.2573 0.5357 + 23.0875%
Overall Mean 0.4792 0.4688 0.4593 0.4834 0.3303 0.3549 0.3025 0.5882 + 13.5327%
Table 4: The performance of VLLMs and humans on the GC@$1$ metric, evaluated using 5 randonly drawn images per sample/image category. the best performance achieved by an MLLM is underlined.
Figure 2 presents a correlation matrix among GC@$T$ scores and the benchmarks mentioned above, where the overall GC@$T$ scores are averaged over the GC@$T$ scores of all MME categories. The strong correlations with the OpenCompass scores incorporating the results of multiple meticulously crafted benchmarks indicate that MMECeption provides a comprehensive evaluation that may complement existing benchmarks. Further, GenCeption appears to effectively measure an MLLM’s tendency to hallucinate, as demonstrated by the strong correlations with HallusionBench. While these observations are further emphasized by the correlations with MMStar and AI2D, the only moderate correlations with MME and SEEDBench provide more nuanced insights. As MME displays these moderate correlations also with the other benchmarks, it can be reasoned that it measures dimensions supplement to those measured by other benchmarks and GenCeption. SEEDBench on the other hand correlates strongly with other benchmarks, but only moderately with GC@$T$ scores. This indicates that SEEDBench measures aspects that are also measured by other benchmarks, but fail to be captured by GenCeption. Future research could focus on identifying these aspects to potentially incorporate them into GenCeption.
Table 3 compares GC@$1$ scores using different image generators, OpenAI’s DALL·E 3 Ramesh et al. (2021) and Google DeepMind’s Imagen2 DeepMind (2023). Independent of image generator used, the rankings remain unchanged, except that on visual-intensive samples only, Claude3-Opus scores equally with GPT-4V. This provides evidence that even though DALL·E 3, GPT-4o, and GPT-4V were developed and trained by OpenAI, neither of OpenAI’s models has an advantage over non-OpenAI MLLMs.
Table 4 shows human performance on a subset of 5 randomly drawn images per category compared to the MLLM performance on the same subset of samples. It can be observed that humans outperform all MLLMs, with especially strong differences in performance for the text-intensive categories. The worst performance, relative to humans, is achieved on the code reasoning and text translation categories, the former containing images of code snippets and the latter of phrases written in simplified Chinese characters. The relatively best performance by MLLMs is achieved on the scene and artwork categories, which contain every-day life photos and popular artworks. This demonstrates that there is still substantial space for performance improvement, and that compared to humans, MLLMs still lack relevant competences. It must be noted that human performance here does not constitute an upper bound in possible scores to achieve, and that future generations of MLLMs may well outperform humans.
3.2 Qualitative Results
We qualitatively inspect our results by visualizing generated images together with their cosine similarity and GC@$T$ scores for two seed images across different categories, as shown in Figure 3. This visualization highlights the correlation between these scores and the visual characteristics of the images relative to the seed image. A key observation is that later iterations show an increased tendency to produce imagery deviating from the seed image, as indicated by lower GC@$T$ scores. This serves as an additional qualitative validation of the GenCeption method and the MMECeption benchmark, as using MLLMs scoring higher on the MMECeption benchmark results in generated images that preserve more information from the seed image. For a wider range of examples across MME image categories and corresponding descriptions from each evaluated VLLM, readers are referred to Appendix A.
Figure 3: Demonstration of GenCeption evaluation procedure: the images generated over 3 GenCeption iterations for several MLLMs. The similarity $s^{(t)}$ scores (to the seed image) are shown on the top of images; GC@$1$ and GC@$3$ scores are printed on the bottom of the first and third image, respectively.
4 Discussion and Future Directions
This study validates the GenCeption method focusing on the visual modality only. Future research is invited to adapt GenCeption to other modalities, such as audio, video, and graphs. For instance, the framework can be initiated with a dataset of audio samples, and MLLMs can iteratively generate and describe the audio content. Similarly, for video and graph data, the process can involve generating textual descriptions of short video clips or graph structures and their recreation. While the core iterative process of GenCeption remains applicable, these extensions require modality-specific generation and embedding models.
The broad skill assessment provided by GenCeption goes along with the limitation that it is difficult to assess which skills contribute most to a high GC@$T$ score. Our analysis indicates that contemporary VLLMs perform poorly on text-intensive tasks while excelling in describing scenes and artworks. Future research could investigate this in a more fine-grained manner by creating datasets requiring specialized skills. For example, datasets could include images of complex emotions, dynamic movements, mechanical processes, or user interfaces. Additionally, combining GenCeption with specifically designed similarity metrics may offer more detailed insights into specific MLLM abilities.
5 Conclusion
In this paper, we introduce GenCeption to enhance the evaluation of rapidly evolving Multimodal Language Models (MLLMs). The GenCeption method addresses key limitations of existing MLLM benchmarks, such as costly data annotation, quick benchmark saturation, training data contamination, leading questions, and the illusion of emergent abilities. Being adaptable to different modalities, the GenCeption method can deliver value as a unified MLLM evaluation method that complements existing MLLM benchmarks.
Our empirical validation using the MMECeption benchmark shows that GenCeption effectively assesses semantic coherence and consistency across modalities, aligning with established VLLM benchmarks. By assessing humans on the MMECeption task, we demonstrate that current VLLMs significantly lag behind human performance, particularly when working with text-intensive images. Future work is encouraged to refine and extend this framework across a wider range of modalities, datasets, and similarity metrics.
Appendix A GenCeption Demonstration
To provide a comprehensive, intuitive and qualitative understanding of the GenCeption procedure and GC@${T}$ metric, we illustrate the input, output, intermediate artifacts, similarity scores, and GC@${T}$ values throughout the GenCeption process. Examples from the visual-intensive and textual-intensive groups are showcased in Figures 5 and 6, respectively. The corresponding seed images and their metadata are presented in Figure 4.
Appendix B Limitations and Societal Impact
The limitations, outlined in Sections 3 and 4, primarily pertain to our initial experimental focus on image-based experiments, excluding other modalities. A critical assumption is the minimal influence of stochastic variability in image generation and MLLM text generation processes. While we have not delved into ethical risks, our framework’s purpose – to assess inter-modality semantic drift and susceptibility to hallucination in MLLMs—is clearly articulated. Societally, the exclusive use of the English language in GenCeption experiments may inadvertently marginalize non-English-speaking user groups.
Appendix C Dataset and Reproducibility
In Sections 1, 2.1, 2.2 and 3 of the main paper, we cite the creators of all artifacts used. Detailed citations can be found in references. The MME dataset is not directly downloadable, and is released for research purposes only upon a request from authors to gain access to it. It does not contain any personally identifying information, as the questions regard visual aspects of the images. We followed the guidelines provided by the authors and respected the intended terms of use. The specific licenses and terms for the use and distribution of publicly available artifacts can be found in the corresponding original papers or GitHub repositories, as cited. As per this research work and aligning with the MME copyrights, we are not releasing this asset. Regarding the created artifacts, we introduce a new metric called GC@$T$, and detail its creation and intended use in Section 2.2 of the main paper. Our study exclusively utilizes images from the MME dataset, omitting textual QA annotations, and generates textual data in the form of English descriptions as part of our methodology. Given the nature of our research centered on quantifying the inter-modality coherence and consistency, we do not apply any data splits. Due to limitations in terms of computational resources, the metrics reported in Table 2 are from a single run.
In our study, we adopt several state-of-the-art models to facilitate our experiments, including Gemini1.5-Pro, GPT-4o, GPT-4V, Claude3-Opus, LLaVa-13B, LLaVa-7B, and mPLUG-Ow12 for text description generation, ViT for image embedding, and DALL·E 3 and Imagen2 for image generation, adhering to default parameter settings as outlined in their original specifications. As we only evaluated existing models but did not train new models, no hyperparamter tuning was applicaple. The text descriptions generated by GPT-4V are obtained through API calls, while experiments involving the other models are conducted on A100 GPUs, totaling approximately 96 GPU hours. Image generation was also performed via a call to OpenAI’s DALL-E 3 API. To compute the GC@$T$ metric, we employ the cosine similarity metric from the Scikit-learn library (Version 1.4.0).
Appendix D Human Annotators
As described in Section 3, we benchmarked the performance of humans at the GC@$T$ task. The 5 human annotators were recruited from the authors’ social circle and were being made aware that they contributed to a research project, with the specific goal of the research project only being disclosed after participation. They were provided with the same instruction prompt as the MLLMs, and given 14 weeks to complete the task. This task does not involve sharing any personal information and the images were carefully evaluated to not be offensive in any way. The participants gave consent that their annotations could be used for scientific research and included in research papers.
Appendix E AI Assistants
AI assistants like GitHub Copilot, ChatGPT, and Perplexity were used to support writing the necessary codebase and find efficient ways to express complex concepts. AI assistant suggestions were always carefully evaluated for correctness and only used after human revision.